i have a few questions here i jut do not understand they are from different sections therefore i am sorry but i could not give it a topic here they are:
what is the greatest monomial factor of 9a-9b?
what is the greatest monomial factor of ax+3x?
what is the greatest monomial factor of d^3-d^2-3d?
factor the polynomial 1/25-x^2
what is not a factor of 3x^3-60+15x^2-12x?
solve each equation using addition or subtraction.(same question here)
5m+n/2=3
4m-3n/9=2
find the values of y in the equation y(y-7)(y-9)=0
solve each equation for all variables values
4y^2=600
36y^2-25=0
(x=3)^2-12x=4
(2x-1)(x=5)=0
z^2=18
if you could show the work that would really help me out.

Answer by stanbon(75887)

what is the greatest monomial factor of 9a-9b?:::::9
-------------------
what is the greatest monomial factor of ax+3x?:::::x
-------------------
what is the greatest monomial factor of d^3-d^2-3d?:::d
--------------------
factor the polynomial 1/25-x^2
Ans: 1/[(5-x)(5+x)]
-----------------------
what is not a factor of 3x^3-60+15x^2-12x?
Ans: Many things are not a factor of that.
----------------------------------------------
solve each equation using addition or subtraction.(same question here)
5m+n/2=3
4m-3n/9=2
Ans: Your post is confusing; use parentheses.
-------------------------------------
find the values of y in the equation y(y-7)(y-9)=0
Ans: y = 0 ; or 7 or 9
-------------------------------------
solve each equation for all variables values
4y^2=600
y^2 = 150
y = 5sqrt(6) or y = -5sqrt(6)
---------------------------------------
36y^2-25=0
(6y-5)(6y+5) = 0
y = 5/6 or y = -5/6
------------------------
(x+3)^2-12x=4
x^2-6x+9 = 4
(x-3)^2-4 = 0
[x-3-2)(x-3+2) = 0
(x-5)(x-1) = 0
x = 5 or x = 1
-------------------------
(2x-1)(x+5)=0
x = 1/2 or x = -5
-------------------------
z^2=18
x^2 - 18 = 0
(x-3sqrt(2))(x+3sqrt(2)) = 0
x = 3sqrt(2) or x = -3sqrt(2)
================================
Cheers,
Stan H.
===========================